• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 16:05:17 on November21,2004

    Entry number 137275

    keyword=Day shift summary

    Shift crew :
    A. Camsonne
    B. Craver
    H. Ibrahim

    Shift summary

    12:22 Two arm run LD2

    12:25 Moving to Optics

    12:30 Optics run

    12:41 Moved back to LD2 back to prodction

    12:48 Bleedthrough 0.15 % ( Hall C at 32 uA now )

    13:22 Rebooted DVCS IOC

    13:30 Restarted CODA on Compton

    15:35 Bleedthrough 1.32%


    Run list

    Two arms
    4886 LD2 1 4.0 203 k TWOARM LD2 15cm
    4887 Optics 1 4.0 202 k TWOARM Optics


    Production
    Run Target Mode Current Event Charge DTime Comment
    4880 LD2 1 4.0 298 k 0.009504 33.78 nDVCS LD2 15cm production
    4881 LD2 1 4.0 298 k 0.009542 35.55 nDVCS LD2 15cm production
    4882 LD2 1 4.0 298 k 0.00953 32.71 nDVCS LD2 15cm production
    4883 LD2 1 4.0 298 k 0.00951 33.62 nDVCS LD2 15cm production
    4884 LD2 1 4.0 298 k 0.00951 37.41 nDVCS LD2 15cm production
    4885 LD2 1 4.0 298 k 0.009521 34.01 nDVCS LD2 15cm production
    4888 LD2 1 4.0 298 k 0.00954 32.22 nDVCS production LD2 15cm
    4889 LD2 1 4.0 298 k 0.009536 32.82 nDVCS production LD2 15cm
    4890 LD2 1 4.0 298 k 0.00953 37.57 nDVCS production LD2 15cm
    4891 LD2 1 4.0 298 k 0.009554 35.39 nDVCS production LD2 15cm


    Compton run list
    9620 75.9 +- 1.1% S/N: 12 no photon detector calibration plot
    9621 75.1 +- 1.6% S/N: 17.3


    Trip list
    8:00-9:00 12 = 5 mins
    9:00-10:00 16 = 6 mins
    10:00-11:00 15 = 3 mins
    11:00-12:00 13 = 6 mins
    12:00-13:00 7 = 2 mins
    13:00-14:00 19 = 6 mins
    14:00-15:00 9 = 3 mins
    15:00-16:00 12 = 4 mins